WRAB operational status:
The station operators restored power with a temporary generator.
IDA engineers visited the station and replaced the backup batteries. Mains power to the station remains down due to issues with the solar power system.
The station is down as of day 2023,048-09:55 due to faulty backup batteries. A station visit to replace the batteries is being planned.
The station is back online after the station operator replaced the batteries.
The station is down due to faulty batteries. IDA engineers are preparing new batteries for a station visit.
The station operator replaced the solar batteries at the station and restored power and data acquisition at 2022,229-17:26.
The station started experiencing several hours of power loss per day on 2022,201. The station operator powered down the station on 2022,207, until new absolyte backup batteries are shipped from UCSD.
The station operator visited the station and replaced the WRAB00 Q330 digitizer.
IDA engineers are onsite and replaced the KS54000 with an STS-6 seismometer, and installed an Episensor strong motion sensor.
There is excessive 1Hz background noise on the vertical channels from the KS54000 seismometer since day 2017,329-15:33 for reasons yet to be determined.
An IDA engineer visited the station and upgraded the power system and replaced the Setra barometer.
Telemetry was down after the station operator replaced the station computers. Telemetry was restored after an IDA engineer reconfigured the connection to the new station computers.
Telemetry to the station is down as of day 2013,043. An IDA engineer has emailed the station operator to investigate.
IDA station WRAB was down from day 2012,254-04:40 to 2012,258-07:14 while an IDA engineer was onsite upgrading the power system.
IDA station WRAB is experiencing daily power failures. An expedition to the station to upgrade the power system is planned.
An IDA engineer visited the station and replaced the STS-2 seismometer with a new Trillium 240 seismometer.
The masses for the STS-2 seismometer drifted out of normal operating range on day 2009,019. The mass position was adjusted and normal operation resumed on day 2009,026.
The station is up as of about day 2008,278 after an IDA engineer visited the station and upgraded the data acquisition system to a Quanterra Q330HR system.
The station is down as of about day 2008,223-15:51 due to problems with the DAS.
The East-West channel (BHE-10) from the STS-2 seismometer has occasional spiking since the DAS was repaired on day 2007,101.
The station was down from about day 2007,275-19:00 to 2007,284-23:57 due to communication problems between the DAS and the ISP.
The short period channels for the STS-2 seismometer were restored as of about day 101-02:18 after the station operator reconfigured the circuit boards in the DAS.
The station is back online as of about day 095-01:01 after the station operator replaced the DAS, recentered the masses for the STS-2 seimometer, and cycled power to the GPS clock. The short period channels for the STS-2 seismometer remain nonseismic for reasons yet to be determined.
The station is down as of about day 346-09:26 due to problems with the DAS. The DAS was shipped back to UCSD for repair(received Jan 3.), and testing of the repaired DAS is in progress.
The wdi-00 channel is restored after the station operator replaced the microbarometer on May 30.
The short period channels from the STS-2 seismometer were spiking from about day 18 (Jan 18) to day 25. The station operator recentered the masses for the STS-2 seismometer and the short period channels returned to normal.
The wdi-00 channel is out due as of 2005,349 (Dec.15) to a faulty microbarometer.
The data acquisition system has been upgraded to a MK8 system. New equipment including a GPS clock, power supplies, radios, and a Paroscientific microbarograph have been installed.
Station operator notes media wear problems, and replaced DAT drive.
Station was down temporarily while maintenance was taking place. Data was not acquired from 2003.07.09 to 2003.07.14.
Clock problem was resolved after it was reset.
Continuous internet connection via satellite link established.
Data streams from the STS2 sensor returned to normal after the masses on the seismometer were recentered by the station operator.
The masses of the STS2 sensors have drifted out of the operating range and will have to be manually readjusted.
New EPROMs installed to correct DSP problem.
By replacing the optical modem on 2002:171-02:30 (June 20, 2002) the timing signal has been fixed.
The station is operating but there are problems with the timing signal being transmitted on the recently-repaired uplink cable. The station configuration is being reviewed to determine how best to bypass the damaged cable.